Missing dimouter, the former platoon Oliver of the rungling brothers eishesy had a kinetic energy is 12000 just primer to hitting the bucket of water. Ik Missy’s mass is 40kg, then what is her speed


1
Expert's answer
2022-01-24T08:45:07-0500

Write the formula for kinetic energy and express speed from it:


"KE=\dfrac 12mv^2→v=\sqrt\dfrac{2E}m=24.5\text{ m/s}."